OpenStudy (anonymous):

Alex drives from X to Y, a distabce of 40km, at an average speed of x km/hr. for the return journey he averages 20km/hr faster. given that the total journey must be completed in an hour and she does not exceed the speed limit of 80km/hr, find x plz plz plz plz help...it has something to do with inequalities

ganeshie8 (ganeshie8):

there is an issue wid this question :- say, she drives at peak speed of 80km/hr for entire journey. then, she wud take half hour for up journey of 40km and half hour for return journey of 40km so going at steady peak speed of 80km/hr she completes the journey in 1 hour. do u see the issue yet ?

OpenStudy (anonymous):

in the wuestion it says she doesnt get higher than 80km/hr...so it would be x<80 but what i want to solve is the minimum speed ..so the range basically..form eg: y<x<80

ganeshie8 (ganeshie8):

there is no range. she has to travel at steady 80kmph if she wants to honor speed limit and time limit. think a bit :)

ganeshie8 (ganeshie8):

speed limit = 80kmph max time limit = 1 hour max

OpenStudy (anonymous):

the return journey is 20 km/hr faster...i want to know..how fast they were b4..so return journey = x+20km/hr... what was the original x value?

ganeshie8 (ganeshie8):

Yes, if u wanto setup inequality it wud look like that

ganeshie8 (ganeshie8):

time inequality :- 40/x + 40/(x+20) <= 1

ganeshie8 (ganeshie8):

for speed u have :- x+20 <= 80 there is no solution for this

OpenStudy (anonymous):

Can you please re-read the original question. There is either a number wrong in translation or it was originally wrong. ganeshie is correct in that there is no solution.

OpenStudy (anonymous):

i rewrote the question word for word....the answer in the book says 71<x<80

OpenStudy (anonymous):

im just going to erase this question from my mind, was trying to do this question for nearly 2 weeks...thanks anyway ;)